Queen Corporation sells on terms of 2/10, 1/30. 70% of customers normally avail of the discounts. Annual sales are P 900,000, 80% of which is made on credit. Cost is approximately 75% of sales.

REQUIRED

a) Average balance of accounts receivable

b) Average investment in accounts receivable.
Answer #1
Sol: Given that ,
Queen corporation sells =2/10,1/30
70% customer avails discounts.
Annual sale= 900000
Credit sales =80%
Cost = 75% of sales.
Average balance of accounts receivables,
Average receivalbles can be computed as sales /receivables turnover.
And receivables turnover = 365/12 months divided by average collection period.
Average collection period ,
Average collection period = (70%*10)*(30%*30) = 16 days., meaning that 70% will avail discount option and remaining 30% wont.
So, average collection period =16 days. Receivables turnover =365/16 = 22.8125.
Average receivables = Sales /turnover,
=900000/22.8125
Average accounts receivables = 39452.05.
Solution b)Let us assume X as average accounts receivable investment

X= 1

year /12 months /Average collection period* Credit sales (where credit sale is 80%)

To compute Average collection period,
Average collection period = (70%*10)*(30%*30) = 16 days.
16 days = (365 * Average accounts receivable investment) / ($900,000 * 80%)
X = (720,000 * 16 ) / 365
X = 31,561.64

Therefore, average accounts receivable investment = 31561.64